A flask containing 300ml of 90% nitrite solution tips over spilling part of its contents. This was subsequently replaced by an equal amount from a mixture of 30% nitrate. If an 80% solution resulted, how much was spilled ?
Answer by stanbon(75887)

A flask containing 300ml of 90% nitrite solution tips over spilling part of its contents. This was subsequently replaced by an equal amount from a mixture of 30% nitrate. If an 80% solution resulted, how much was spilled ?
-------
Equation:
nitrite - nitrite + nitrite = nitrite
-----
0.90*300 - 0.90x + 0.30x = 0.80*300
----------------
90*300 - 90x + 30x = 80*300
-60x = -10*300
x = 50 ml (amt that was spilled)
===========
Cheers,
Stan H.
===========